OpenStudy (anonymous):

need help with 6th grade work!! coin is tossed and a spinner with four equal sections numbered 1 through 4 is spun. What is the probability that the coin will land with heads up and the spinner will stop on 3? A.1/8 B. 1/6 C. 1/3 D. 3/4

OpenStudy (anonymous):

You multiply the probability of the first event by the probability of the second event. In this case it would be (1/2) * (1/4) = 1/8

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

NO Problem at all :) wc If you don't get it feel free to ask me

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

can you help me with a another?

OpenStudy (anonymous):

yes sure!

OpenStudy (anonymous):

A bag contains 5 black marbles and 4 white marbles. A spinner has 3 equal sections labeled 1 through 3. What is the probability of drawing a black marble from the bag and spinning a 1 on the spinner? A. 5/9 B. 1/2 C. 2/9 D. 5/27

OpenStudy (anonymous):

i think its C

OpenStudy (anonymous):

How many marbles are in the bag total?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

8

OpenStudy (anonymous):

4+5=??

OpenStudy (anonymous):

How many man i will give you steps but you need to do math.

OpenStudy (anonymous):

And 5 + 4 is not eight

OpenStudy (anonymous):

9*

OpenStudy (anonymous):

i click the wrong one

OpenStudy (anonymous):

Ok now what is the probability of pulling a black marble

OpenStudy (anonymous):

5

OpenStudy (anonymous):

wait its 5/9 ?

OpenStudy (anonymous):

5 what? Is it certain you will pull the marble five times?

OpenStudy (anonymous):

Yes 5/9

OpenStudy (anonymous):

Now whats the probability of spinning a 1

OpenStudy (anonymous):

yes! thanks!

OpenStudy (anonymous):

Now 5/9 is not the answer its the probability of pulling a black!!!

OpenStudy (anonymous):

Whats the probability of spinning a 1?

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

fiya

OpenStudy (anonymous):

Fiya?

OpenStudy (anonymous):

i got 40 percent

OpenStudy (anonymous):

mi go 40% >:0

OpenStudy (anonymous):

How many times the #1 appears/total numbers